OpenStudy (anonymous):

NEED HELP FAST Find the product of (2x + 5y)2. 4x2 + 20xy + 25y2 4x2 − 20xy + 25y2 4x2 − 25y2 4x2 + 25y2

OpenStudy (sirm3d):

this expansion is applicable \[\large (a+b)^2=(a)^2+2(a)(b)+(b)^2\] where \[\large a = 2x; \space b=5y\]

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

i was thinking it was D but a and b confused me